OpenStudy (anonymous):

find the derivative of y with respect to x, t or theta as appropriate. y= ln(cos(ln theta))

OpenStudy (anonymous):

its very simple

OpenStudy (anonymous):

please help me

OpenStudy (anonymous):

-(1/cos(ln theta))*(sin(ln theta)*1/theta

OpenStudy (stamp):

\[\frac{d(ln(cos(ln(\theta)))}{dx}\]

OpenStudy (stamp):

Should be dTheta

OpenStudy (anonymous):

the answer is - (tan (ln theta))/ theta

OpenStudy (anonymous):

i just dont no how to get to that

OpenStudy (stamp):

\[\frac{1}{cos(ln(\theta))}*-sin(ln(\theta))*\frac{1}{\theta}\]\[\frac{-sin(ln(\theta))}{cos(ln(\theta))}\frac{1}{\theta}\]\[=\frac{-tan(ln(\theta))}{\theta}\]Let me know if we need to go over anything.

OpenStudy (anonymous):

thank you this explains it :)

OpenStudy (stamp):

Ok good luck in calculus

OpenStudy (anonymous):

thank you its really hard. just trying to keep up. have a test coming up i just hope to pass it. thank you for ur help!

OpenStudy (stamp):

Sure thing. Private message me if you ever want to discuss more problems.

OpenStudy (anonymous):

ok thank you!